OpenStudy (anonymous):

Solve 8x2 + 7x - 15 = - 7.

OpenStudy (anonymous):

say 8x2 you mean 8x^2, so for that move the -7 from the left side making the equation 8x^2 + 7x - 8 = 0 From there you simply use the quadratic equation: x = 1/16 (-7-sqrt(305)) x = 1/16 (sqrt(305)-7)

OpenStudy (anonymous):

What would the answer come out to. Would it be something like \[\frac{ -7+\sqrt{305} }{ 16 }\]

OpenStudy (anonymous):

@Littlemsadventuretime

OpenStudy (anonymous):

yes exactly:)

OpenStudy (anonymous):

Is that the only answer or is there another

OpenStudy (anonymous):

Nope thats it. I mean thats what i got

OpenStudy (anonymous):

Ok great thank you!

OpenStudy (anonymous):

Really quick would the answer to -x2 - 3x + 1 = 0 be \[\frac{ -3+\sqrt{13} }{ 2 }\]

OpenStudy (anonymous):

@Littlemsadventuretime

OpenStudy (anonymous):

yes:)